OpenStudy (anonymous):

calculate the no of atoms in 52 moles of helium

OpenStudy (unklerhaukus):

Avogadro's number is the number of particles in one mole of any substance

OpenStudy (chmvijay):

1mole of He= 6.023*10^23 atoms of He 52 moles of He = 2* 6.023*10^23

OpenStudy (whpalmer4):

@chmvijay that should be \(52 * 6.023*10^{23} = 3.13196*10^{25} \) not \(2*6.023*10^{23} = 1.2046*10^{24}\)

OpenStudy (chmvijay):

yaaa ur right :) 2 shouldn't be there :P

OpenStudy (chmvijay):

5 is missed there :P